OpenStudy (anonymous):

In the coordinate plane, which equation describes the locus of points 5 units from the point (–2, 2)?

jimthompson5910 (jim_thompson5910):

hint: if you keep coming back to something over and over again, then you are going in _____

OpenStudy (anonymous):

sane

jimthompson5910 (jim_thompson5910):

lol no

jimthompson5910 (jim_thompson5910):

in circles

OpenStudy (anonymous):

close enough lol

jimthompson5910 (jim_thompson5910):

so that's what the locus they want you find, the equation of the circle

jimthompson5910 (jim_thompson5910):

in this case, center: (-2,2) radius: 5

OpenStudy (anonymous):

this is confusing their answers are A. (x + 2)2 + (y – 2)2 = 25 B. (x – 2)2 + (y + 2)2 = 5 C. (x – 2)2 + (y + 2)2 = 25 D. (x + 2)2 + (y – 2)2 = 5

jimthompson5910 (jim_thompson5910):

remember the last problem how I used (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

well all you do is plug in h = -2, k = 2 and r = 5 and simplify things a bit

jimthompson5910 (jim_thompson5910):

(x-h)^2+(y-k)^2=r^2 (x-(-2))^2+(y-2)^2=5^2 I'll let you handle the last step

jimthompson5910 (jim_thompson5910):

able to figure it out?

jimthompson5910 (jim_thompson5910):

or no?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

(x-(-2))^2+(y-2)^2=5^2 (x+2)^2+(y-2)^2=25 don't worry if you don't get it the first few times, keep practicing though

OpenStudy (anonymous):

you are my favorite person in the whole world right now, thanks! lol

jimthompson5910 (jim_thompson5910):

lol aw i feel special, glad I could be of help